OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

(y-y1)=m(x-x1) m=slope (x1,y1) is the point simplify to get answer

OpenStudy (anonymous):

here slope is 3 since product of perpendicular slopes is -1

OpenStudy (anonymous):

the slope of perpendicular line is the negative reciprocal? +1/3 m (y+1)=1/3(x-6) y + 1 = 1/3 x -2 y= 1/3 -3